# KKL Observer Synthesis for Nonlinear Systems via Physics-Informed Learning

### v1

@@ 4.4 Learning Algorithm
In our decoupled learning procedure, we first learn $\hat{\theta}^{\star}$ in (15) for the forward map $\hat{\mathcal{T}}_{\theta}:\mathcal{X}\rightarrow\mathcal{Z}$ by fitting the dataset $(X,Z)$ using (14) as a loss function. The learned forward map is then used to generate a new dataset $(Z^{\prime},X^{\prime})$, with sampled initial conditions
$(\hat{\mathcal{T}}_{\hat{\theta}^{\star}}(x_{0}^{1}),x_{0}^{1}),\dots,(\hat{\mathcal{T}}_{\hat{\theta}^{\star}}(x_{0}^{p}),x_{0}^{p}).$
The change in order between $Z^{\prime}$ and $X^{\prime}$ in the dataset notation means that $Z^{\prime}$ serves in this instance as the features and $X^{\prime}$ the labels. Our access to the forward map $\hat{\mathcal{T}}_{\hat{\theta}^{\star}}:\mathcal{X}\to\mathcal{Z}$ allows us to be independent of the truncation process and directly map the sampled initial conditions in $\mathcal{X}$ to their values in $\mathcal{Z}$, therefore allowing us to choose the distribution of the initial conditions while avoiding system blow-ups in backward time. The parameters $\hat{\eta}^{\star}$ in (17) for the inverse map $\hat{\mathcal{T}}_{\eta}^{*}$ is then learned by minimizing (16) on the new dataset $(Z^{\prime},X^{\prime})$. The act of decoupling the training of $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$ therefore avoids the risk of conflicting gradients between (14) and (16), while ensuring that $\hat{\mathcal{T}}_{\eta}^{*}$ is indeed trained using data from the correct distribution.
@@ 5 Robustness of Learned KKL Observer
The neural networks $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$ are mere approximations of $\mathcal{T}$ and $\mathcal{T}^{*}$, respectively. The approximation errors will influence the estimation performance of the observer if we use the neural networks instead of the original transformation maps. However, only the approximation error associated with the inverse $\hat{\mathcal{T}}_{\eta}^{*}$ explicitly affects the state estimation error, and the approximation error associated with $\hat{\mathcal{T}}_{\theta}$ only implicitly affects the state estimation error as is apparent in the proof of Proposition 16 below.
For any $z\in\mathcal{Z}$, the true inverse $\mathcal{T}^{*}(z)$ can be written as
$\mathcal{T}^{*}(z)=\hat{\mathcal{T}}_{\eta}^{*}(z)+\mathcal{E}^{*}(z)$ (18)
where $\mathcal{E}^{*}(z)$ is the approximation error of the learned inverse map $\hat{\mathcal{T}}_{\eta}^{*}$ at $z\in\mathcal{Z}$. Using (18), the KKL observer (7) can be written as
$\begin{array}[]{ccl}\dot{\hat{z}}(t)&=&A\hat{z}(t)+By(t),\quad\hat{z}(0)=\hat{z}_{0}\\ \hat{x}(t)&=&\hat{\mathcal{T}}_{\eta}^{*}(\hat{z}(t))+\mathcal{E}^{*}(\hat{z}(t))\end{array}$ (19)
where the approximation error $\mathcal{E}^{*}(\hat{z}(t))$ is an unknown signal corrupting the state estimate $\hat{x}(t)$. Moreover, the model (1) of the system is never perfect in real-world applications, and there are underlying uncertainties that could also influence the state estimation error. In this section, we provide robustness guarantees in terms of explicit input-to-state stability bounds for the estimation error under both the approximation error and the system uncertainties.
@@ 5.1 Problem Setup and Assumptions
In practice, there are uncertainties in system dynamics, i.e.,
@@
where $w(t)\in\mathbb{R}^{n_{x}}$ and $v(t)\in\mathbb{R}^{n_{y}}$ denote the model uncertainty and the measurement noise, respectively. In the presence of uncertainties, the objective of the observer changes from steering the estimation error to zero as in (4) to steering it “close” to zero. That is, the observer must ensure that the estimation error (3) satisfies an input-to-state stability bound
$\|\xi(t)\|\leq\beta(\|\xi_{0}\|,t)+\gamma_{1}(\|w_{[0,t]}\|_{L^{\infty}})+\gamma_{2}(\|v_{[0,t]}\|_{L^{\infty}})$ (21)
where $\xi(t)$ is the estimation error defined in (3), $\xi_{0}=\xi(0)$ is the initial estimation error, $\beta$ is a class-$\mathcal{KL}$ function [^5] A function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is of class $\mathcal{KL}$ if for each fixed $t\in\mathbb{R}_{\geq 0}$, $\beta(\cdot,t)$ is of class $\mathcal{K}$, and for each fixed $r\in\mathbb{R}_{\geq 0}$, $\beta(r,t)\to 0$ as $t\to\infty$. that asymptotically converges to zero as $t\to\infty$, and $\gamma_{1},\gamma_{2}$ are class-$\mathcal{K}_{\infty}$ functions. From (21), it is evident that in the absence of uncertainty $w(t)$ and noise $v(t)$, the estimate $\hat{x}(t)$ asymptotically converges to the true state $x(t)$, thus satisfying (4). Moreover, (21) imposes robustness criterion on the observer’s state estimate under uncertainties. In other words, as the magnitudes of the uncertainties increase, (21) ensures that there is a graceful degradation of the state estimate, so the estimation error remains bounded as long as the uncertainties are bounded.
Let $x(t;x_{0},w)$ denote the state of (20a) at time $t\in\mathbb{R}_{\geq 0}$ initialized at $x(0)=x_{0}$ and driven by $w_{[0,t]}$.
@@
    Boundedness: $w\in L^{\infty}(\mathbb{R};\mathbb{R}^{n_{x}})$ and $v\in L^{\infty}(\mathbb{R};\mathbb{R}^{n_{y}})$ with $\|w\|_{L^{\infty}}<\overline{w}$ and $\|v\|_{L^{\infty}}<\overline{v}$, for some $\overline{w},\overline{v}>0$.
@@
    $\|x(t;x_{0},w)-x(t;x_{0},0)\|\leq\psi(\|w_{[0,t]}\|_{L^{\infty}}).$
Assumption 10(i) is a standard assumption in robust state estimation [15]. Assumption 10(ii) requires that the deterministic system (1) does a good job in describing the uncertain system (20). Providing such a guarantee is widely studied in system identification, and the reader is referred to [39, 37, 33, 1] for more details.
@@
where
$\sigma(t)=h(\bar{x}(t))-h(x(t))-v(t)$ (23)
with $x(t)\coloneqq x(t;x_{0},w)$ and $\bar{x}(t)\coloneqq x(t;x_{0},0)$.
Note that $\sigma(t)\in\mathbb{R}^{n_{y}}$ in (23) remains bounded for bounded uncertainty $w(t)$ and noise $v(t)$ due to Assumption 10. That is, for every $x(t),\bar{x}(t)\in\mathcal{X}$ and $t\in\mathbb{R}_{\geq 0}$,
@@
where $\overline{w},\overline{v}$ are given in Assumption 10(i) and $\ell_{h}<\infty$ with
@@
$\|v(t)\|\leq\sqrt{n_{y}}\|v(t)\|_{\infty}\leq\sqrt{n_{y}}\|v_{[0,t]}\|_{L^{\infty}}\leq\sqrt{n_{y}}\overline{v}.$

### v2

@@ 5.4 Remarks on the Learning Method
In our sequential procedure, we first learn $\theta$ for the forward map $\hat{\mathcal{T}}_{\theta}:\mathcal{X}\rightarrow\mathcal{Z}$ by solving (17) on the dataset $S_{1}$. The learned forward map is then used to generate a new dataset $S_{2}$. The parameters $\eta$ for the inverse map $\hat{\mathcal{T}}_{\eta}^{*}$ are then learned by solving (19) on the new dataset $S_{2}$. Decoupling the training of $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$ therefore avoids the risk of conflicting gradients between (16) and (18), while ensuring that $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$ are indeed trained using data from correct distributions.
The formulation of the empirical physics-informed risk $\hat{R}_{S_{1}}(\theta)$ in (16) constitutes a semi-supervised learning framework. The data-fit term $\|z-\hat{\mathcal{T}}_{\theta}(x)\|^{2}$ is computed on the labeled data $S_{\mathrm{data}}$. Conversely, the physics-informed risk term $\|\mathcal{P}_{\theta}(x)\|^{2}$ is evaluated on the unlabeled collocation data $S_{\mathrm{pde}}$. This semi-supervised structure allows us to enforce the physical constraint (6) over a large, densely-sampled portion of the state space ($N_{\mathrm{pde}}\gg N_{\mathrm{data}}$), which acts as a powerful data-agnostic regularizer and prevents overfitting to the labeled $S_{\mathrm{data}}$.
6 Non-Asymptotic Learning Guarantees
The learning method proposed in Section 5 computes the approximations $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$. A critical question remains: does this learning procedure provide any guarantee that the resulting maps are “good” approximations of the true, unknown maps $\mathcal{T}$ and $\mathcal{T}^{*}$? The practical utility of the learned observer depends on this approximation quality. If the learned inverse map $\hat{\mathcal{T}}_{\eta}^{*}$ is a poor approximation of the true $\mathcal{T}^{*}$, the resulting estimate $\hat{x}(t)=\hat{\mathcal{T}}_{\eta}^{*}(\hat{z}(t))$ will be inaccurate, irrespective of the observer’s dynamic properties.
This section provides theoretical guarantees for our learning-based design of KKL observers. We establish that the approximation error of the learning procedure is bounded. This error is quantified by the inverse map error $R_{\mathcal{T}^{*}}(\eta)$ defined in (14). Proving that $R_{\mathcal{T}^{*}}$ can be made small provides the foundation for our learning method.
6.1 Lipschitz Assumption
Given that the neural network’s activation functions are Lipschitz continuous, the neural network map is also Lipschitz continuous. Thus, it is natural to make the following assumption.
Assumption 10.
The neural network $\hat{\mathcal{T}}_{\eta}^{*}$ is Lipschitz continuous over $\mathcal{Z}$, where $\mathcal{Z}\supseteq\mathcal{T}(\mathcal{X})$. In particular, for every $z,\hat{z}\in\mathcal{Z}$, there exists $\ell_{\eta}\in\mathbb{R}_{>0}$ such that
$\|\hat{\mathcal{T}}_{\eta}^{*}(z)-\hat{\mathcal{T}}_{\eta}^{*}(\hat{z})\|\leq\ell_{\eta}\|z-\hat{z}\|$. (20)
Estimating a neural network’s Lipschitz constant is a topic of interest in the machine learning community [48, 52, 23, 26]. For feedforward ReLU networks, [20] shows that the local Lipschitz constant estimation problem can be reduced to a semidefinite program.
6.2 Generalization Guarantees
Recall that our learning method is a sequential procedure. First, we learn $\hat{\mathcal{T}}_{\theta}\in\mathcal{H}_{\theta}$ by minimizing the empirical physics-informed risk $\hat{R}_{S_{1}}(\theta)$. Then, we learn $\hat{\mathcal{T}}_{\eta}^{*}\in\mathcal{H}_{\eta}$ by minimizing the empirical reconstruction risk $\hat{R}_{S_{2}}(\eta\mid\theta)$. The inverse map error $R_{\mathcal{T}^{*}}$ is thus affected by the outcomes of both stages. In the following lemma, we relate $R_{\mathcal{T}^{*}}$ to the reconstruction risk $R_{2}(\eta\mid\theta)$ in (15), which is empirically minimized in the second stage, and the forward map error $R_{\mathcal{T}}(\theta)$ in (11), which is propagated from the first stage.
Lemma 11.
Let Assumption 10 hold. Then, the true inverse error $R_{\mathcal{T}^{*}}(\eta)$ in (14) is bounded by
$R_{\mathcal{T}^{*}}(\eta)\leq 2R_{2}(\eta\mid\theta)+2\ell_{\eta}^{2}R_{\mathcal{T}}(\theta)$ (21)
where $R_{2}(\eta\mid\theta)$ is the reconstruction risk (15), $R_{\mathcal{T}}(\theta)$ is the forward map error (11), and $\ell_{\eta}$ is the Lipschitz constant in (20).
{pf}
Let $\|\cdot\|_{L_{2}(\mathcal{D}_{\mathcal{X}})}$ denote the $L_{2}$ norm with respect to the data distribution $\mathcal{D}_{\mathcal{X}}$. Then,
$\sqrt{R_{\mathcal{T}^{*}}}$ $=\|\mathcal{T}^{*}(\mathcal{T})-\hat{\mathcal{T}}_{\eta}^{*}(\mathcal{T})\|_{L_{2}(\mathcal{D}_{\mathcal{X}})}$
$\leq\|\mathcal{T}^{*}(\mathcal{T})-\hat{\mathcal{T}}_{\eta}^{*}(\hat{\mathcal{T}}_{\theta})\|_{L_{2}(\mathcal{D}_{\mathcal{X}})}$
$\hskip 60.00009pt+\|\hat{\mathcal{T}}_{\eta}^{*}(\hat{\mathcal{T}}_{\theta})-\hat{\mathcal{T}}_{\eta}^{*}(\mathcal{T})\|_{L_{2}(\mathcal{D}_{\mathcal{X}})}$
$\leq\sqrt{R_{2}(\eta\mid\theta)}+\ell_{\eta}\|(\hat{\mathcal{T}}_{\theta}-\mathcal{T})\|_{L_{2}(\mathcal{D}_{\mathcal{X}})}$
$=\sqrt{R_{2}(\eta\mid\theta)}+\ell_{\eta}\sqrt{R_{\mathcal{T}}}$
where we used $x=\mathcal{T}^{*}(\mathcal{T}(x))$, the triangle inequality, Assumption 10, and definitions (15) and (11). Squaring both sides using the inequality $(a+b)^{2}\leq 2a^{2}+2b^{2}$ gives the desired result (21). ∎
Lemma 11 states that the inverse map error $R_{\mathcal{T}^{*}}$ is bounded by the reconstruction risk $R_{2}(\eta\mid\theta)$ and the forward error $R_{\mathcal{T}}$, which is amplified by the Lipschitz constant of the learned inverse map. We now bound these two terms using the generalization bound presented in Section 4.
Lemma 12.
Let $\mathcal{H}_{\eta}$ be the hypothesis class for the learned inverse map $\hat{\mathcal{T}}_{\eta}^{*}$. Let $\mathcal{G}_{\eta}=\{(x^{\prime},z^{\prime})\mapsto\|x^{\prime}-\hat{\mathcal{T}}_{\eta}^{*}(z^{\prime})\|^{2}:\hat{\mathcal{T}}_{\eta}^{*}\in\mathcal{H}_{\eta}\}$ be the associated loss class, with finite pseudo-dimension $d_{\eta}=\mathrm{Pdim}(\mathcal{G}_{\eta})$, where $z^{\prime}=\hat{\mathcal{T}}_{\theta}(x)$. Let $N_{2}$ be the number of data points in the dataset $S_{2}=\{(x_{j}^{\prime},z_{j}^{\prime})\}$. Assume the empirical reconstruction risk $\hat{R}_{S_{2}}(\eta\mid\theta)$ is bounded by $M_{\eta}$. Then, for any $\delta>0$, with probability at least $1-\delta$, we have
$R_{2}(\eta\mid\theta)\leq\inf_{\hat{\eta}\in\mathcal{H}_{\eta}}\hat{R}_{S_{2}}(\hat{\eta}\mid\theta)+2\mathcal{C}_{\eta}(N_{2},\mathcal{H}_{\eta},\delta)$ (22)
where $\mathcal{C}_{\eta}(N_{2},\mathcal{H}_{\eta},\delta)=M_{\eta}\sqrt{\frac{2d_{\eta}\ln(eN_{2}/d_{\eta})}{N_{2}}}+M_{\eta}\sqrt{\frac{\ln(1/\delta)}{2N_{2}}}$, $R_{2}$ defined in (15), and $\hat{R}_{S_{2}}$ in (18).
Lemma 13.
Let $\mathcal{H}_{\theta}$ be the hypothesis class for the learned forward map $\hat{\mathcal{T}}_{\theta}$. Let $\mathcal{G}_{\theta}$ be the associated loss class for the empirical physics-informed risk $\hat{R}_{S_{1}}(\theta)$ in (16), with finite pseudo-dimension $d_{\theta}=\mathrm{Pdim}(\mathcal{G}_{\theta})$. Let $N_{1}$ be the number of data points in $S_{1}$. Assume $\hat{R}_{S_{1}}(\theta)\leq M_{\theta}$. Then, for any $\delta>0$, with probability at least $1-\delta$,
$R_{1}(\theta)\leq\inf_{\theta\in\mathcal{H}_{\hat{\theta}}}\hat{R}_{S_{1}}(\hat{\theta})+2\mathcal{C}_{\theta}(N_{1},\mathcal{H}_{\theta},\delta)$ (23)
where $\mathcal{C}_{\theta}(N_{1},\mathcal{H}_{\theta},\delta)=M_{\theta}\sqrt{\frac{2d_{\theta}\ln(eN_{1}/d_{\theta})}{N_{1}}}+M_{\theta}\sqrt{\frac{\ln(1/\delta)}{2N_{1}}}$, $R_{1}$ defined in (13), and $\hat{R}_{S_{1}}$ in (16).
Moreover, since $R_{\mathcal{T}}(\theta)\leq R_{1}(\theta)$, the right-hand side of (23) also provides a high-probability bound on $R_{\mathcal{T}}$.
Lemmas 12 and 13 are direct applications of the generalization bound in Lemma 9 to the empirical risk minimization problems (19) and (17). In these lemmas, the assumptions that the empirical risks are bounded by constants $M_{\eta}$ and $M_{\theta}$ are justified by the compactness of the state space $\mathcal{X}$ (Assumption 2). Since the neural networks $\hat{\mathcal{T}}_{\theta},\hat{\mathcal{T}}_{\eta}^{*}$ utilize continuous activation functions and the system maps $f,h$ and the transformation $\mathcal{T}$ are continuous, the empirical risk functions, which are composed of squared errors and PDE residuals, are continuous functions defined on a compact set. Therefore, by the Extreme Value Theorem, they strictly admit finite upper bounds.
6.3 Main Theorem and Discussion
We provide a high-probability bound on the true $L_{2}$ approximation error $R_{\mathcal{T}^{*}}$ of the learned inverse map $\hat{\mathcal{T}}_{\eta}^{*}$ by combining the results of the above lemmas.
Theorem 14.
Let $\hat{\mathcal{T}}_{\theta}\in\mathcal{H}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}\in\mathcal{H}_{\eta}$ be the maps learned by the proposed learning method. Let Assumption 10 hold. Then, for any $\delta>0$, with probability at least $1-2\delta$, it holds that
$R_{\mathcal{T}^{*}}(\eta)\leq 2\left(\hat{R}_{2}(\eta\mid\theta)+2\mathcal{C}_{\eta}(N_{2},\mathcal{H}_{\eta},\delta)\right)\\ +2\ell_{\eta}^{2}\left(\hat{R}_{1}(\theta)+2\mathcal{C}_{\theta}(N_{1},\mathcal{H}_{\theta},\delta)\right)$ (24)
where $R_{\mathcal{T}^{*}}(\eta)$ is the inverse map error in (14), $\hat{R}_{2}(\eta)=\inf_{\hat{\eta}\in\mathcal{H}_{\eta}}\hat{R}_{S_{2}}(\eta\mid\theta)$ is the approximation error of the inverse map, $\hat{R}_{1}=\inf_{\hat{\theta}\in\mathcal{H}_{\theta}}\hat{R}_{S_{1}}(\theta)$ is the approximation error of the forward map, and $\mathcal{C}_{\eta}$ and $\mathcal{C}_{\theta}$ are the explicit complexity terms defined in Lemmas 12 and 13, which vanish as the data sizes $N_{2}$ and $N_{1}$ tend to infinity.
The proof follows from Lemmas 11, 12, and 13. Below, we discuss some implications of the theorem and present guidelines for the learning algorithm.
Theorem 14 provides an explicit, non-asymptotic bound on the inverse map error $R_{\mathcal{T}^{*}}$. It provides a theoretical guarantee for the learning algorithm presented in Section 5. It demonstrates that $R_{\mathcal{T}^{*}}$ can be made small, provided the empirical approximation errors, $\hat{R}_{2}$ and $\hat{R}_{S_{1}}$, and the estimation errors, $\mathcal{C}_{\theta}$ and $\mathcal{C}_{\eta}$, are small. This is achieved by using sufficiently rich network architectures $\mathcal{H}_{\theta}$ and $\mathcal{H}_{\eta}$ (e.g., deep and wide) to be able to represent the true maps $\mathcal{T}$ and $\mathcal{T}^{*}$, leveraging the universal approximation property of neural networks. Moreover, the estimation error terms $\mathcal{C}_{\eta}$ and $\mathcal{C}_{\theta}$ decrease as the data sizes $N_{1}$ and $N_{2}$ increase. This justifies the data-generation procedure and shows that a larger dataset results in a lower inverse map error.
The bound (24) also depends on the forward map error $R_{\mathcal{T}}$, which represents the error from learning the forward map $\hat{\mathcal{T}}_{\theta}$ and can be bounded as $R_{\mathcal{T}}\leq R_{1}\leq\hat{R}_{1}+2\mathcal{C}_{\theta}$ (Lemma 13). However, this error $R_{\mathcal{T}}$ is amplified by $\ell_{\eta}^{2}$ in (24). Since the Lipschitz constant $\ell_{\eta}$ depends on the norm of the neural network weights, regularizing the weights of the inverse map $\hat{\mathcal{T}}_{\eta}^{*}$ results in a smaller $\ell_{\eta}$.
The role of our physics-informed learning method is as follows. The bound (24) on $R_{\mathcal{T}^{*}}$ depends on $R_{\mathcal{T}}$, which is bounded by $R_{1}$. Therefore, minimizing the PDE residual $\|\mathcal{P}_{\theta}\|^{2}$ (the physics-informed component of $\hat{R}_{S_{1}}(\theta)$ in (16)) is not just a heuristic; it is essential for provably reducing the forward map error $R_{\mathcal{T}}$ and, consequently, for minimizing the inverse map error $R_{\mathcal{T}^{*}}$. The guarantee that we can learn an accurate inverse map $\hat{\mathcal{T}}_{\eta}^{*}$ is the fundamental prerequisite for analyzing the observer’s performance in the presence of noise and uncertainties, which we will show in the next section.
Additionally, the physics-informed risk component ($\hat{R}_{1}$ and $\mathcal{C}_{\theta}$) in the bound (24) highlights the theoretical benefit of our approach. A standard supervised method (i.e., with $\nu=0$) would search for a minimizer within the entire, unconstrained hypothesis class $\mathcal{H}_{\theta}$. In contrast, incorporating the PDE constraint acts as a powerful inductive bias. It implicitly restricts the optimization to an effective hypothesis space, $\mathcal{H}_{\theta,\mathrm{PDE}}\subset\mathcal{H}_{\theta}$, which contains only those functions that are physically plausible (i.e., that nearly satisfy the KKL PDE (6)). This effective space is significantly less complex than $\mathcal{H}_{\theta}$, implying that its pseudo-dimension $d_{\theta,\mathrm{PDE}}$ is much smaller than the pseudo-dimension $d_{\theta}$ of the unconstrained class. Consequently, the complexity term $\mathcal{C}_{\theta}$ (defined in Lemma 13) is substantially reduced. This demonstrates that the physics-informed approach not only finds a more accurate solution (lower $\hat{R}_{1}$) but also achieves a better generalization guarantee (lower estimation error) from a finite dataset $S_{1}$.
@@ 7 Robustness of Learned KKL Observer
The neural networks $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$ aproximate $\mathcal{T}$ and $\mathcal{T}^{*}$, respectively. Approximation errors affect the observer’s estimation performance when we use neural networks instead of the original transformation maps. For any $z\in\mathcal{Z}$, the true inverse $\mathcal{T}^{*}(z)$ can be written as
$\mathcal{T}^{*}(z)=\hat{\mathcal{T}}_{\eta}^{*}(z)+\mathcal{E}_{\eta}^{*}(z)$ (25)
where $\mathcal{E}_{\eta}^{*}$ denotes the approximation error of the learned inverse map $\hat{\mathcal{T}}_{\eta}^{*}$. The learned KKL observer is given by
$\begin{array}[]{ccl}\dot{\hat{z}}(t)&=&A\hat{z}(t)+By(t),\quad\hat{z}(0)=\hat{z}_{0}\\ \hat{x}(t)&=&\hat{\mathcal{T}}_{\eta}^{*}(\hat{z}(t)).\end{array}$ (26)
The estimate obtained by this observer is affected by the approximation error $\mathcal{E}_{\eta}^{*}(\hat{z}(t))$, which acts as an unknown noise. Additionally, the system model (1) is rarely perfect in practice, and underlying uncertainties can affect the estimation performance. In this section, we provide robustness guarantees for the estimation error under both approximation error and system uncertainties.
@@ 7.1 Problem Setup and Assumptions
Consider an uncertain nonlinear system
@@
where $w(t)\in\mathbb{R}^{n_{x}}$ and $v(t)\in\mathbb{R}^{n_{y}}$ denote the model uncertainty and the measurement noise, respectively. For uncertain systems, the objective of the observer changes from steering the estimation error to zero as in (4) to steering it “close” to zero. That is, the observer must ensure that the estimation error (3) satisfies an input-to-state stability bound
$\|\xi(t)\|\leq\beta(\|\xi_{0}\|,t)+\gamma_{1}(\|w_{[0,t]}\|_{L_{\infty}})+\gamma_{2}(\|v_{[0,t]}\|_{L_{\infty}})$ (28)
where $\xi(t)$ is the estimation error defined in (3), $\xi_{0}=\xi(0)$ is the initial estimation error, $\beta$ is a class-$\mathcal{KL}$ function [^5] A function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is of class $\mathcal{KL}$ if (i) for each fixed $t\in\mathbb{R}_{\geq 0}$, $\beta(r,t)$ is of class $\mathcal{K}$, and (ii) for each fixed $r\in\mathbb{R}_{\geq 0}$, $\beta(r,t)$ is decreasing in $t$ and is such that $\beta(r,t)\to 0$ as $t\to\infty$. that asymptotically converges to zero as $t\to\infty$, and $\gamma_{1},\gamma_{2}$ are class-$\mathcal{K}_{\infty}$ functions. From (28), it is evident that in the absence of uncertainty $w(t)$ and noise $v(t)$, the estimate $\hat{x}(t)$ asymptotically converges to the true state $x(t)$, thus satisfying (4). Moreover, (28) imposes a robustness criterion on the observer’s state estimate under uncertainties. In other words, as the magnitudes of the uncertainties increase, (28) ensures that there is a graceful degradation of the state estimate, so the estimation error remains bounded as long as the uncertainties are bounded.
Let $x(t;x_{0},w)$ denote the state of (27a) at time $t\in\mathbb{R}_{\geq 0}$ initialized at $x(0)=x_{0}$ and driven by uncertainty $w_{[0,t]}$. When $w\equiv 0$, the state trajectory is denoted by $x(t;x_{0})$.
@@
    Boundedness: It holds that $\|w\|_{L_{\infty}}<\overline{w}$ and $\|v\|_{L_{\infty}}<\overline{v}$, for some known $\overline{w},\overline{v}>0$.
@@
    $\|x(t;x_{0},w)-x(t;x_{0})\|\leq\psi(\|w_{[0,t]}\|_{L_{\infty}}).$
Assumption 15(i) is quite standard in robust state estimation [17]. Assumption 15(ii) requires that the deterministic system (1) does a good job in describing the uncertain system (27). Providing such a guarantee is widely studied in system identification, and the reader is referred to [37, 35, 31, 1] for more details.
@@
where $\sigma(t)=h(\bar{x}(t))-h(x(t))-v(t)$ with $x(t)\coloneqq x(t;x_{0},w)$ and $\bar{x}(t)\coloneqq x(t;x_{0})$. Note that $\sigma(t)\in\mathbb{R}^{n_{y}}$ remains bounded due to Assumption 15. That is, for every $x(t),\bar{x}(t)\in\mathcal{X}$ and $t\in\mathbb{R}_{\geq 0}$,
@@
where $\overline{w},\overline{v}$ are from Assumption 15(i) and $\ell_{h}<\infty$ with
@@
$\|v(t)\|\leq\sqrt{n_{y}}\|v(t)\|_{\infty}\leq\sqrt{n_{y}}\|v_{[0,t]}\|_{L_{\infty}}\leq\sqrt{n_{y}}\overline{v}.$